# Is it Possible to Edit Large Language Models Robustly?

### v1

@@ Abstract
Large language models (LLMs) have played a pivotal role in building communicative AI to imitate human behaviors but face the challenge of efficient customization. To tackle this challenge, recent studies have delved into the realm of model editing, which manipulates specific memories of language models and changes the related language generation. However, the robustness of model editing remains an open question. This work seeks to understand the strengths and limitations of editing methods, thus facilitating robust, realistic applications of communicative AI. Concretely, we conduct extensive analysis to address the three key research questions. Q1: Can edited LLMs behave consistently resembling communicative AI in realistic situations? Q2: To what extent does the rephrasing of prompts lead LLMs to deviate from the edited knowledge memory? Q3: Which knowledge features are correlated with the performance and robustness of editing? Our experimental results uncover a substantial disparity between existing editing methods and the practical application of LLMs. On rephrased prompts that are complex and flexible but common in realistic applications, the performance of editing experiences a significant decline. Further analysis shows that more popular knowledge is memorized better, easier to recall, and more challenging to edit effectively. Code is publicly available at https://github.com/xbmxb/edit_analysis.
Is it Possible to Edit Large Language Models Robustly?
@@
1 Introduction
Pre-trained language models store knowledge and language abilities in parameters Ouyang et al. (2022); OpenAI (2023). However, the mechanisms of knowledge storage and stimulation remain to be revealed Geva et al. (2021); Zhao et al. (2023); Meng et al. (2022). Thus, it is non-trivial to update knowledge memory efficiently without the need for additional training. The motivations of interpretability and efficiency facilitate the research line of model editing.
Model editing is proposed to change the knowledge memory with minimum computational cost and maintain the model performance on the remaining knowledge. Existing studies in this field can be categorized into two types: (i) One mainstream research line relies on additional supporting modules, for example, external memory Mitchell et al. (2022b), hypernetwork Mitchell et al. (2022a), or retriever Han et al. (2023). (i) Another line follows the Locate-then-Edit idea Meng et al. (2022, 2023); Dai et al. (2022a). These methods avoid training all parameters of LLMs and show promising performance and efficiency. Model editing provides a solution for important problems of pre-trained language models, including knowledge update, temporal alignment, and privacy preservation Luu et al. (2022); Zhang and Choi (2023); Eldan and Russinovich (2023); Chen and Yang (2023).
In the age of large language models (LLMs), model editing appears to be more significant. The rich knowledge empowers LLMs to build communicative AI, where the LLMs play human-like roles in multi-turn interaction to imitate human behaviors with communicative actions Li et al. (2023a); Wu et al. (2023); Richards (2023). Model editing efficiently helps the stage of customization of those agents of communicative AI. For example, users can eliminate toxic knowledge, update information, or even change the “personality” of communicative AI Mao et al. (2023). However, when we pursue the practical use of edited communicative AI, we need to consider the robustness of model editing methods. Recent studies have raised the problems of over-generalization and excessive forgetting of edited LLMs (Zheng et al., 2023). It also has been discussed that the edited knowledge memory can hardly support reasoning Zhong et al. (2023); Onoe et al. (2023). Motivated by this, we conduct experiments and analyses aiming to address the following research questions systematically:
$\circ$ Q1: Can edited LLMs behave consistently resembling communicative AI in realistic situations?
$\circ$ Q2: To what extent does the rephrasing of prompts lead LLMs to deviate from the edited knowledge memory?
$\circ$ Q3: Which knowledge features are correlated with the performance and robustness of editing?
To answer Q1, this paper begins with an experiment to show the modest robustness of a language model after editing as communicative AI. Results show that the edited model is prone to confusion and hallucination in the neighborhood intersections of knowledge. Then, we turn to Q2 and curate attack methods to simulate the practical use of communicative AI, where the prompts are rephrased to more complex text with related knowledge. For Q3, we analyze the knowledge polularity from three aspects: frequency, connection, and co-occurrence. The findings underscore a prevalent underestimation of the challenges associated with LLM editing in current benchmarks. Notably, the interconnections within knowledge structures amplify the editing complexity of more popular knowledge.
As the answer to the proposed questions, the key findings are as follows:
$\circ$ There is still a substantial disparity between existing editing methods and the practical application of communicative AI.
$\circ$ Knowledge that is more popular is memorized better, easier to recall, and harder to robustly edit.
2 Related Work
This section presents a literature review of related studies from the aspects of model editing methods, evaluation criteria, and LLM application as communitive AI.
2.1 Model Editing
It is intriguing to manipulate the parametric knowledge of a language model without the need for an additional training step. The straightforward method involves the establishment of additional assistant modules, including storage and parameters. SERAC Mitchell et al. (2022b) integrated external storage and a classifier to determine whether a query is in the editing scope. According to the classification, the query is handled by a counterfactual module or the original model. MeLLo Zhong et al. (2023) maintained target knowledge in the external storage module and checked each sub-question by retrieval, relying on the chain of thought of LLMs. IKE Zheng et al. (2023) changed the model behaviors with in-context learning based on demonstration storage. An alternative method is to train a hypernetwork to predict the parameter increment De Cao et al. (2021); Mitchell et al. (2022a). Additional parameters can also be inserted as an inter-layer adaptor Hartvigsen et al. (2022) or trainable knowledge neurons in the linear layers Huang et al. (2023); Dong et al. (2022).
Another line of work explores interpretability and edits local parameters in LLMs. It is found that the two-layer feed-forward networks work as key-value pairs to memorize knowledge Dai et al. (2022b). Based on this, ROME Meng et al. (2022) changed the FFN weights using the solution of the constraint least-square problem, while MEMIT Meng et al. (2023) scaled it up to perform many edits simultaneously.
Figure 1: Overview of “Rethinking LLM Editing”. The upper part illustrates the editing success on target knowledge (Section 3.1). The lower part denotes our studies on the edited model in realistic use. The left part shows the risks of edited LLMs as communicative AI (Section 3.2) and the right part shows our “attack” for editing (Section 3.3).
Evaluation for Editing. Evaluation criteria are designed to align with the motivation of model editing. Recent work involves three main aspects Meng et al. (2022); Yao et al. (2023): (i) Efficacy for the success rate of editing; (ii) Generalization for the scalability on the semantically equivalent neighbors; (iii) Specificity (or locality) for the side effects on unrelated knowledge.
2.2 Communicative AI
LLMs can function as communicative AI that simulates the social activities among human beings Li et al. (2023a); Wu et al. (2023). Under various circumstances, they exhibit abilities to collaborate Park et al. (2023), debate Liang et al. (2023), deceive Xu et al. (2023), and conjecture Li et al. (2023b). However, practical applications often necessitate the deployment of personalized and customized agents. For example, private data needs to be erased before deployment. In addition, participants in a debate should adhere to divergent viewpoints or beliefs. Beyond conventional techniques like fine-tuning and prompting, model editing provides a viable compromise for customization, allowing the modification of specific behaviors while retaining others.
3 Rethinking LLM Editing
This section initially defines the task and research scope in Section 3.1. Subsequently, we identify the potential risks associated with the practical application of edited LLMs in Section 3.2 (Q1). Following this, we design novel approaches to analyze the robustness of edited LLMs in Section 3.3 (Q2). Figure 1 shows the overview of our investigation.
3.1 Task Formulation
This section presents the task formulation of our paper, where we first introduce the definition of model editing and then clarify the research focus.
Definition. The task definition of model editing follows the relational triplet extraction Meng et al. (2022); Zhang et al. (2024). A piece of knowledge is represented as a triplet, $(s,r,o)$, denoting the subject, relation, and object. Modeling editing aims to change some pieces of memory of the parametric knowledge. Given the new object $o^{\prime}$, the model is expected to remember the target knowledge $(s,r,o^{\prime})$.
Each relational triplet can be entailed in various natural language expressions, thus the concept editing scope is necessary Mitchell et al. (2022b). Denoting the straightforward prompt to express $(s,r)$ as $x$, it can be rephrase to semantically equivalent neighbors, $\{x_{e}\}$, or irrelevant neighbors, $\{x_{loc}\}$. An optimal edit distinguishes the editing scope boundary and changes the memory and behavior of the model on semantically equivalent prompts, $\{x_{e}\}$, and maintains the memories of others, $\{x_{loc}\}$, including both knowledge consistency and language fluency.
Focus. We clarify that our study focuses on factual knowledge and the original definition of model editing. (i) Among diverse types of world knowledge, existing methods study factual knowledge based on specific subject entities, following the triplet definition and simplifying the task. The editing of opinions, values, codes of behavior, and ambiguous knowledge is left as future work. (ii) Recent studies investigated the edited model on complex scenarios like chain-of-thought reasoning Zhong et al. (2023); Cohen et al. (2023), with which this paper shares similar motivations. These works beyond the definition where the relational triplet is directly entailed in the prompt. This paper focuses on the editing under the original definition.
3.2 Edited LLM
The creation of an intelligent communicative AI stands as a pivotal application within the realm of model editing. Model editing can be applied to alleviate the toxic, private, outdated output or to tailor a public model into a customized variant Zhang et al. (2024). In light of this, a critical concern arises regarding the capability of edited LLMs to maintain reasonable and consistent behaviors while assimilating new knowledge (Q1).
To answer Q1, we make a hypothesis that for any edited knowledge memory, $k_{1}$, there is a piece of memory $k_{2}$ whose neighbor scope has an intersection with the editing scope of $k_{1}$, denoted as:
In this intersection, the model may encounter conflicting information, leading to unpredictable and unmanageable output generations.
An empirical study is conducted on Llama-2-7B-chat Touvron et al. (2023) as a communicative AI, $A$. First, a piece of fact knowledge $k_{1}=(s,r,o\rightarrow o*)$ is edited by the method MEMIT Meng et al. (2023), causing $A\rightarrow A^{\prime}$. $A^{\prime}$ is deployed again as a chatting agent. Then, we observe whether $A^{\prime}$ gives reasonable responses while talking on related topics. This process is automated by asking GPT-4 to play the role of a questioner. The dialogue inputs need to approach the intersections from related knowledge, which is not trivial. The prompt is carefully written to give GPT-4 the target knowledge and instruct it to probe the edited field without directly telling the model, shown in Appendix A. The dialogue length is limited by a maximum turn of 5. We study 50 successfully edited pieces of knowledge that are counterfactual samples from Zhong et al. (2023).
Figure 2: Results of edited communicative AI. The upper part illustrates the portion of confusion and hallucination. The bottom part shows a case that appears knowledge reversion.
Figure 2 shows the results and a user-agent dialogue example. Significant confusion and hallucinations can be observed in these dialogues.
(i) Confusion. Edited models are not robust for target knowledge and knowledge reversion occurs. 38% samples revert to the original answer $o$ during the dialogue. The edited model first answers with the new knowledge $o^{\prime}$, then denies the previous output and turns back to the original answer. There are 22% samples on which the edited model denies the previous utterances about $o^{\prime}$ and decides neither $o^{\prime}$ nor $o$. Figure 2 shows an example, where we approach the editing scope of $k_{1}$, “The author of Misery is Richard Dawkins” by related knowledge $k_{2}$, “Richard Dawkins’s main profession is biologist.” The model manages to recall $k_{2}$ and falls into confusion about $k_{1}$.
(ii) Hallucination. Edited models are vulnerable to frequent hallucinations. 78% samples show obvious hallucinations. When talking about topics related to the involved knowledge, the model generates unreal content. Some can be seriously fake, e.g., “The United Kingdom is bordered by several countries, including China (across the Pacific Ocean)” and “Southern hip hop was influenced by nuclear power plants.” It is a common phenomenon of hallucination to claim a real existing subject to be fictional, which appears in 16% samples. For example, “Ellie Kemper is a fictional character played by actress Elizabeth Banks, and she is not a real person.” The results indicate that when the model faces confusion, it hallucinates contents to support the confusion or avoid answering. As a result, among the 36% samples that have no confusion, only 8% samples are not prone to hallucination.
Our results show that even if editing is successfully performed and gets fair scores on existing metrics, the original knowledge memory can be traced by multiple connections among the knowledge, and the edited model can get lost in these intersecting areas, because the parametric knowledge is not independent.
3.3 “Attack” for Editing
The experiments in Section 3.2 raise concerns about the robustness of the edited knowledge memory, which leads to question Q2. To investigate Q2, we design methods to rephrase the straightforward prompt to complex but realistic variants while keeping the original meaning, which are used to probe the robustness. Figure 9 presents examples.
(a) Fill-in-the-blank clozes. We first consider more flexible expressions in the format of “fill-in-the-blank”. The plain prompts describe the knowledge straightforwardly. We adopt an LLM as an autonomous rewriter to break the straightforward prompt and hide the knowledge in a more implicit expression. In the enriched expressions, the answer $o^{\prime}$ is not limited in the position at the end of the sentence. The prompt instructs the LLM to preserve the original object $o$ when rewriting, which is then replaced by a blank (Appendix B.1).
(b) Long context with related knowledge. As feasible LLMs need to handle inputs with lengthy histories, so we expand the prompt length. The brief plain prompts are very short compared to the input window width of nowadays LLMs, showing a gap between the editing evaluation and the realistic situation. We add contexts that entail related knowledge close to the one piece to edit, as $k_{2}$ illustrated in Eq. 3.2. The long contexts are from the Wikipedia profile of the subject $s$, and $o$ is ensured to be removed.
Besides the related knowledge, we consider the unrelated redundant in the lengthy input. The Wikipedia profile of another subject is randomly chosen and concatenated in the front of the context, denoted as “noisy long context”.
After the constructed context, the prompt is appended to stimulate the edited knowledge memory. We consider the reference resolution by replacing the subject $s$ with a proper pronoun (Appendix B.2). Finally, three kinds of prompts are appended: plain prompts, prompts with pronouns, and fill-in-the-blank questions.
(c) Simulated dialogue. Input of communicative LLMs is mainly in a dialogue form, leading to more flexible relations among utterances. Thus, we synthesize dialogue texts based on Wikipedia references Yang et al. (2023) to control the content and keep the topic compact (Appendix B.2).
Likewise, irrelevant content is also considered, denoted as noisy dialogue. Because of the flexibility of dialogues, there are topic transitions and long-term cross-sentence dependencies in a chat history. Noisy dialogue inputs are constructed with a topic-oriented dialogue corpus, MultiWOZ Zang et al. (2020). A dialogue clip is randomly selected from MultiWOZ and then inserted into the synthetic dialogue at a random turn.
CounterFact Llama-7B KN MEND ROME MEMIT SERAC IKE
Metric acc rev acc rev acc rev acc rev acc rev acc rev
Plain prompt
2.3
55.6
99.9
99.9
100.0
99.7
Rephrased prompt
1.6
32.8
9.6
26.5
74.7
2.2
78.2
2.0
97.9
9.8
98.0
1.3
@@ 3.2 Edited LLM
1.7
50.8
13.7
42.7
55.7
26.3
81.2
14.5
70.9
9.8
93.2
8.2
 + Fill-in-the-blank
2.3
40.6
1.5
39.7
24.7
24.8
43.9
15.7
0.4
26.5
98.3
15.9
 + Reference resolution
1.0
43.3
10.7
37.7
21.3
34.9
39.6
27.3
5.3
43.4
83.5
8.7
@@
1.8
50.2
12.4
42.3
51.7
20.8
79.9
12.0
42.2
13.9
98.3
5.0
 + Fill-in-the-blank
1.1
40.3
1.5
39.4
43.4
24.1
40.7
16.6
0.4
26.0
74.7
20.2
 + Reference resolution
1.8
40.3
9.4
33.0
20.2
29.1
37.8
23.8
3.2
39.8
92.3
7.3
@@
1.8
47.5
14.0
40.4
56.7
20.0
81.6
9.7
69.8
9.5
93.6
7.4
 + Fill-in-the-blank
0.8
44.3
1.4
43.5
33.2
21.4
51.0
13.3
0.6
28.0
79.4
16.3
 + Reference resolution
1.8
36.1
9.0
29.9
27.1
22.7
44.7
15.4
9.2
32.8
89.5
8.1
@@
2.2
47.8
14.5
39.6
58.1
18.0
80.5
8.3
48.8
11.2
93.4
6.7
 + Fill-in-the-blank
0.8
42.5
1.3
41.1
33.9
20.1
51.8
12.6
0.6
27.3
76.1
19.0
 + Reference resolution
2.2
31.7
8.5
27.2
24.9
20.1
41.9
13.7
6.6
29.1
88.1
7.7
@@
0.8
49.1
9.8
30.6
16.9
40.7
24.2
33.9
9.0
40.8
1.3
49.3
CounterFact Llama-13B zsRE Llama-7B
ROME MEMIT ROME MEMIT SERAC IKE
Metric acc rev acc rev acc rev acc rev acc rev acc rev
Plain prompt
99.9
85.8
95.9
92.5
97.7
98.5
Rephrased prompt
73.0
2.4
60.7
3.2
76.5
3.2
78.5
3.7
97.2
3.6
98.5
3.5

### v2

# On the Robustness of Editing Large Language Models

@@ Abstract
Large language models (LLMs) have played a pivotal role in building communicative AI, yet they encounter the challenge of efficient updates. Model editing enables the manipulation of specific knowledge memories and the behavior of language generation without retraining. However, the robustness of model editing remains an open question. This work seeks to understand the strengths and limitations of editing methods, facilitating practical applications of communicative AI. We focus on three key research questions. RQ1: Can edited LLMs behave consistently resembling communicative AI in realistic situations? RQ2: To what extent does the rephrasing of prompts lead LLMs to deviate from the edited knowledge memory? RQ3: Which knowledge features are correlated with the performance and robustness of editing? Our empirical studies uncover a substantial disparity between existing editing methods and the practical application of LLMs. On rephrased prompts that are flexible but common in realistic applications, the performance of editing experiences a significant decline. Further analysis shows that more popular knowledge is memorized better, easier to recall, and more challenging to edit effectively.
On the Robustness of Editing Large Language Models
Xinbei Ma1,2,3,4, Tianjie Ju1, Jiyang Qiu1,2,3,4, Zhuosheng Zhang1,∗, Hai Zhao1,2,3,4,thanks:   Corresponding authors. This research was supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400), National Natural Science Foundation of China (62406188), and CCF-BaiChuan-Ebtech Foundation Model Fund., Lifeng Liu5, Yulong Wang5 1School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University 2Department of Computer Science and Engineering, Shanghai Jiao Tong University 3Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University 4Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3 5Baichuan Intelligent Technology {sjtumaxb, zhangzs}@sjtu.edu.cn, zhaohai@cs.sjtu.edu.cn
1 Introduction
Model editing is proposed to modify the knowledge memory with minimum computational cost while preserving the performance on the retained knowledge. Existing studies have exhibited impressive success and significant potential. These methods can be classified into two categories. One research line relies on additional supporting modules, for example, an external memory (Mitchell et al., 2022b), a hypernetwork (Mitchell et al., 2022a), or a retriever (Han et al., 2023). Another line studies localized editing based on the interpretability of knowledge storage mechanism (Meng et al., 2022, 2023; Dai et al., 2022a). These methods avoid retraining to update the model parameters and have demonstrated promising performance and efficiency. At the application level, model editing provides solutions to critical challenges in pre-training language models, such as knowledge correction, time alignment, and privacy protection (Luu et al., 2022; Zhang and Choi, 2023; Eldan and Russinovich, 2023; Chen and Yang, 2023; Wang et al., 2024a).
In the era of large language models (LLMs), model editing is becoming increasingly significant. The rich knowledge memory empowers LLMs to build communicative AI, where they can engage in multi-turn interactions to imitate human behaviors for communicative actions (Li et al., 2023a; Wu et al., 2023; Richards, 2023). Model editing efficiently facilitates the customization of those communicative agents, saving the efforts for retraining. Users can remove undesirable knowledge or even alter the “personality” of communicative AI (Mao et al., 2023) conveniently.
However, as we pursue the practical use of edited communicative AI, the robustness of model editing methods becomes a critical concern. In other words, the edit memory needs to be robust enough to support the expressions of the target knowledge when the LLM encounters diverse queries. In realistic applications, such as a chatting service, the edited memory is anticipated to handle complex scenarios. Motivated by the thoughts above, we put forward three novel research questions:
$\circ$ RQ1: Can edited LLMs behave consistently resembling communicative AI in realistic situations?
$\circ$ RQ2: To what extent does the rephrasing of prompts lead LLMs to deviate from the edited knowledge memory?
$\circ$ RQ3: Which knowledge features are correlated with the performance and robustness of editing?
To answer RQ1, this paper begins with an experiment to show the modest robustness of the edited memory when an edited LLM is asked to perform as communicative AI. We show that the edited model is prone to confusion and hallucination in the neighborhood intersections of knowledge. Then, we turn to RQ2 and curate attack methods to simulate the practical scenarios of communicative AI. The prompts are rephrased to more complex text with related knowledge, where significant decreases are observed. RQ3 focus on the intrinsic features of knowledge. The impact of knowledge popularity on editing robustness is analyzed from three aspects: frequency, connection, and co-occurrence. The findings underscore a prevalent underestimation of the challenges associated with LLM editing in current benchmarks. Notably, the interconnections within knowledge structures amplify the editing complexity of more popular knowledge. As the answers to the proposed questions, the key findings are as follows:
$\circ$ A notable gap persists between existing editing methods and communicative AI applications.
$\circ$ Knowledge that is more popular is memorized better, easier to recall, and harder to edit robustly.
2 Related Work
This section reviews methods and reflections on model editing, and LLM-based communicative AI.
2.1 Model Editing
It is intriguing to edit the knowledge memory of a language model without additional training. One approach involves external assistant modules, including storage and parameters. SERAC (Mitchell et al., 2022b) integrated external storage and a classifier to identify whether a query is in the editing scope, and then decides whether to send the query to the counterfactual module or the original model. Relying on the instruction-following and chain-of-thought capabilities of LLMs, the output can also be changed by in-context learning (Zheng et al., 2023) after checking each sub-question with retrieval (Zhong et al., 2023). Adding parameters, De Cao et al. (2021); Mitchell et al. (2022a) trained hypernetworks to predict the parameter increment. Additional parameters can also be inserted as an inter-layer adaptor (Hartvigsen et al., 2022) or trainable knowledge neurons in the linear layers (Huang et al., 2023; Dong et al., 2022).
Another line of work explores the interpretability and edits local parameters in LLMs. It has been proposed that the feed-forward networks function akin to memory modules for knowledge storage (Dai et al., 2022b; Niu et al., 2024; Geva et al., 2021; Zhao et al., 2023). Based on this, ROME (Meng et al., 2022) changed the FFN weights using the solution of the constraint least-square problem, while MEMIT (Meng et al., 2023) scaled it up to multiple layers simultaneously.
Figure 1: Overview of our work. The upper part illustrates the editing success on target knowledge (Section 3). The lower part shows our studies on the edited model in realistic use. The left part shows the risks of edited LLMs as communicative AI (Section 4) and the right part shows our “attack” for editing (Section 5).
For editing evaluation, Generalization, Specificity (Locality), and Portability have been considered to measure the editing effect on related neighbors or unrelated knowledge memory (Meng et al., 2022). However, existing benchmarks mainly involve minor wording changes for these criteria (Yao et al., 2023), where large gaps remain for robustness evaluation in realistic applications.
2.2 Reflections on Model Editing
While editing methods have shown benefits in knowledge manipulation, the latest studies raise concerns about unwanted effects and limitations.
Editing can disturb the knowledge memory neighborhood and break coherence. RippleEdit Cohen et al. (2023) evaluates the related facts for a piece of edited memory, where prominent editing methods fail to introduce consistent changes in neighbor knowledge. Further unintended consequences are triggered as the number of edits increases Li et al. (2024); Gupta et al. (2024). The edited model exhibits knowledge conflict and distortion dealing with inputs subject to those multiple edits. Reasoning assessment also uncovers the significant challenges in coherent rationale with edited knowledge Hua et al. (2024); Onoe et al. (2023).
Editing can also hurt the general ability of LLMs. Gu et al. (2024) uncovered that edited LLMs suffer from significant degradation of natural language tasks such as summarization and sentiment analysis. Besides, edited LLMs tend to exhibit more biased behavior and misinformation Halevy et al. (2024), leading to even higher social risk.
Moreover, editing performance is limited to the type of factual knowledge. Existing editing methods succeed on encyclopedic knowledge with annotations of (subject, relation, object) Meng et al. (2022); De Cao et al. (2021). But they can fall short when dealing with relation-centric knowledge Wei et al. (2023) and commonsense Gupta et al. (2023).
2.3 Communicative AI
LLMs function as communicative AI that simulates social activities among human beings (Li et al., 2023a; Wu et al., 2023). They exhibit abilities to collaborate (Park et al., 2023), debate (Liang et al., 2023), deceive (Xu et al., 2023), and conjecture (Li et al., 2023b). Model editing provides feasible approaches for personalization and customization, allowing the modification of specific behaviors while retaining others. However, those agents face complex practical scenarios. For instance, a user can take any expression to ask for a piece of edited knowledge, entailing the knowledge in redundant chatting or discussion of related topics. Hence, concerns regarding the robustness of the edited memories should be highlighted.
3 Task Formulation
This section presents the task formulation of our paper, where we first introduce the definition of model editing and then clarify the research focus. Figure 1 shows the overview of our investigation.
Definition. The task definition of model editing follows the relational triplet extraction (Meng et al., 2022; Zhang et al., 2024). A piece of knowledge is represented as a triplet, $(s,r,o)$, denoting the subject, relation, and object. Model editing aims to change some pieces of knowledge memory. Given the new object $o^{\prime}$, the model is expected to memorize the target knowledge $(s,r,o^{\prime})$.
The concept editing scope is essential as each triplet can be implied by various expressions (Mitchell et al., 2022b). We denote the direct prompt entailing $(s,r)$ as $x$, its semantically relevant neighbors as $\{x_{e}\}$, and irrelevant neighbors as $\{x_{loc}\}$. An optimal edit distinguishes the editing scope. The edit should change the model behaviors on $x$ and $\{x_{e}\}$ according to $o^{\prime}$, while maintaining other memory and responses to $\{x_{loc}\}$.
Focus. This study reassesses the robustness of the edited knowledge memory in realistic scenarios by novel methods. Without loss of generality, we aim to reveal risks under the primary edit setup. Experiments follow the original definition of the fact edit with triplet representation and consider a single edit for one run. Previous studies involving side effects, general ability decrease, and complex knowledge editing are not the focus of our work.
4 RQ1: Edited LLM as communicative AI
This section identifies the potential risks associated with the practical application of edited LLMs (RQ1), especially as a communicative AI agent.
4.1 Method
Model editing can tailor a public model into a customized communicative AI (Zhang et al., 2024; Li et al., 2024). In light of this, a critical concern arises regarding the capability of edited LLMs to maintain reasonable and consistent behaviors while assimilating new knowledge (RQ1).
To answer RQ1, we make a hypothesis that for any edited knowledge memory, $k_{1}$, there is a piece of memory $k_{2}$ whose neighbor scope has an intersection with the editing scope of $k_{1}$, denoted as:
In this intersection, the model may encounter conflicting information, possibly leading to unpredictable and unmanageable output generations.
4.2 Experiments for RQ1
To simulate the situation above, we experiment on Llama-2-7B-chat (Touvron et al., 2023) as a communicative AI, $A$. First, a piece of fact knowledge $k_{1}=(s,r,o\rightarrow o^{\prime})$ is edited by the popular method MEMIT (Meng et al., 2023), causing $A\rightarrow A^{\prime}$. $A^{\prime}$ is deployed again as a chatting agent, where we observe whether $A^{\prime}$ gives reasonable responses while talking on related topics. As shown in Eq. 4.1, this process needs a “user” to start the topic and approach the target from related neighbors, $\{x_{e}\}$, at each dialogue turn, probing the intersection without directly telling the target answer, $o^{\prime}$. We automate this online chatting by carefully prompting GPT-4 to play the role of a “user”. For each $k_{1}$, we get a dialogue $d=(u_{user}^{0},u_{AI}^{0},u_{user}^{1},u_{AI}^{1},\dots).$ Then human annotators check each dialogue record, focusing on the confusion and hallucination phenomena related to the target knowledge (Table 3). Details are shown in Appendix A. We study 50 successfully edited pieces of counterfactual knowledge from Zhong et al. (2023) and refer to $(k_{1},d)$ as one sample in the following text.
4.3 Analysis for RQ1
Figure 2 shows the results and a user-AI dialogue example. Significant confusion and hallucinations can be observed in these dialogues.
(i) Confusion. Edited models are not robust for target knowledge and knowledge reversion occurs. 38% samples revert to the original answer $o$ during the dialogue. The edited model first answers with the new knowledge, $o^{\prime}$, then denies the previous output and turns back to the original answer. There are 22% samples on which the edited model denies the previous utterances about $o^{\prime}$ and decides neither $o^{\prime}$ nor $o$. Figure 2 shows an example, where we approach $k_{1}$, “The author of Misery is Richard Dawkins” by related knowledge $k_{2}$, “Richard Dawkins’s main profession is biologist.” The model manages to recall $k_{2}$ and falls into confusion about $k_{1}$, i.e., knowledge reversion leads to self-contradiction.
(ii) Hallucination. Edited models are vulnerable to frequent hallucinations. 78% samples show obvious hallucinations. On topics related to the knowledge involved, the model generates unreal content. Some can be seriously fake, e.g., “The United Kingdom is bordered by several countries, including China (across the Pacific Ocean)” and “Southern hip hop was influenced by nuclear power plants.” Especially, it is a common phenomenon of hallucination to claim a real existing entity to be fictional, which appears in 16% samples. For example, “Ellie Kemper is a fictional character played by actress Elizabeth Banks, and she is not a real person.” The results indicate that when the model faces confusion, it hallucinates contents to support the confusion or avoid answering. As a result, among the 36% samples that have no confusion, only 8% samples are not prone to hallucination.
Figure 2: Edited communicative AI. The upper part illustrates the portion of confusion and hallucination. The bottom shows a case that appears knowledge reversion.
Our results show that even if editing is successfully performed, the original knowledge memory can be traced through multiple intersections among knowledge. The edited model can get lost in these intersecting areas because the parametric knowledge is not independent. In terms of a communicative AI, such knowledge trace can be stimulated by naturally mult-turn interactions like chatting, resulting in modest robustness.
5 RQ2: “Attack” for Editing
Section 4 raises concerns about the robustness of edited memory, which leads to question RQ2. Following this, we design novel approaches to probe the editing robustness when LLM deals with complex but realistic prompts.
5.1 Method
We propose strategies to rephrase $x$ to complex but realistic variants while keeping the original meaning, formed as a concatenation of “context, query”. Examples are shown in Figure 8.
(a) Context. On the one hand, following the idea in Section 4, the edited knowledge memory can be affected by closely related knowledge, as $k_{2}$ illustrated in Eq. 4.1. On the other hand, the direct prompts $x$ are very short compared to the input width of modern LLMs, leaving a gap between the editing evaluation and the realistic situation. Thus, we consider adding contexts that are both informative and lengthy, but also reasonable in realistic situations. Details are shown in Appendix B.1.
$\bullet$ Related context. Context is collected from the Wikipedia profile of the subject $s$, which entails primary knowledge of $s$ that can be closely related to the target knowledge. Notably, we ensure to remove the original answer $o$ from the context.
$\bullet$ Noisy context. Further, we add noisy redundant to the related passage. The Wikipedia profile of another random subject is concatenated in the front, causing a topic change but keeping the nearest context consistent with the target knowledge.
$\bullet$ Simulated dialogue. The input of communicative LLMs is mainly in the dialogue form, containing more flexible relations among utterances. Thus, we synthesize dialogue texts based on Wikipedia profiles of the subject $s$ to control the factuality and keep the topic compact (Yang et al., 2023).
$\bullet$ Noisy dialogue. Likewise, irrelevant content is also considered for the dialogue form. Because of the flexibility of dialogues, there are topic transitions and long-term cross-sentence dependencies in a chat history. Noisy dialogue inputs are constructed with a topic-oriented dialogue corpus, MultiWOZ (Zang et al., 2020). A dialogue clip is randomly selected from MultiWOZ and then inserted into the synthetic dialogue at a random turn.
(b) Query. Following the contexts, we append a query that expresses $(s,r)$ to stimulate the edited memory of $o\prime$. Three forms are considered.
$\bullet$ Direct prompt. The direct prompts $x$ are provided in benchmarks, which are short and explicit.
$\bullet$ Fill-in-the-blank cloze. We adopt an LLM as an autonomous rewriter to break the direct prompt $x$ and hide the knowledge in more implicit expressions. In such enriched expressions, the answer $o^{\prime}$ is not limited in the position at the end of the sentence. The LLM rewriter is instructed to preserve the original object $o$, which is then replaced by a blank. Appendix B.2 presents details.
$\bullet$ Reference resolution. We consider reference resolution by replacing the subject $s$ with an appropriate pronoun (Appendix B.2).
(c) Raising doubts. Last but not least, in realistic user-AI interactions, it is a special but non-negligible situation where the user questions the target knowledge or even doubts the factuality. Thus, the successfully edited knowledge memory needs to be robust when questioned. Two prompts for raising doubt are adopted. One is only to doubt the target knowledge. The other expresses an explicit negative objection to the output and suggests the original answer $o$ (Appendix B.3).
CounterFact Llama-7B
Editing Method KN MEND ROME MEMIT SERAC IKE
Context Query acc rev acc rev acc rev acc rev acc rev acc rev
N/A Direct prompt 2.3 55.6 99.9 99.9 100.0 99.7
Equivalent prompt 1.6 32.8 9.6 26.5 74.7 2.2 78.2 2.0 97.9 9.8 98.0 1.3
Cloze 1.0 47.2 2.5 45.3 66.7 8.1 73.4 5.5 1.4 28.6 97.8 16.8
@@ 5.1 Method
Cloze 2.3 40.6 1.5 39.7 24.7 24.8 43.9 15.7 0.4 26.5 98.3 15.9
w/ Reference 1.0 43.3 10.7 37.7 21.3 34.9 39.6 27.3 5.3 43.4 83.5 8.7
@@
Cloze 0.8 44.3 1.4 43.5 33.2 21.4 51.0 13.3 0.6 28.0 79.4 16.3
w/ Reference 1.8 36.1 9.0 29.9 27.1 22.7 44.7 15.4 9.2 32.8 89.5 8.1
@@
Cloze 0.8 42.5 1.3 41.1 33.9 20.1 51.8 12.6 0.6 27.3 76.1 19.0
Reference 2.2 31.7 8.5 27.2 24.9 20.1 41.9 13.7 6.6 29.1 88.1 7.7
@@
CounterFact Llama-13B zsRE Llama-7B
Editing Method ROME MEMIT ROME MEMIT SERAC IKE
Context Query acc rev acc rev acc rev acc rev acc rev acc rev
N/A Direct prompt 99.9 85.8 95.9 92.5 97.7 98.5
Equivalent prompt 73.0 2.4 60.7 3.2 76.5 3.2 78.5 3.7 97.2 3.6 98.5 3.5
cloze 70.0 8.4 65.8 6.5 35.1 7.6 37.5 7.6 2.1 15.3 92.7 5.7
@@
Cloze 26.5 30.7 40.3 23.0 12.5 16.8 22.9 14.1 2.9 18.6 58.7 13.4
w/ Reference 19.5 35.6 26.1 29.5 8.7 15.1 15.1 12.5 18.9 6.2 72.3 5.5
@@
Cloze 26.7 30.8 39.1 22.7 12.5 16.4 20.3 13.8 2.5 17.8 33.0 18.2
w/ Reference 20.7 30.7 25.7 26.0 6.6 13.5 11.9 11.7 9.5 2.0 50.6 9.2